# Nano Drone-based Indoor Crime Scene Analysis*

### v1

@@ Abstract
Technologies such as robotics, Artificial Intelligence (AI), and Computer Vision (CV) can be applied to crime scene analysis (CSA) to help protect lives, facilitate justice, and deter crime, but an overview of the tasks that can be automated has been lacking. Here we follow a speculate prototyping approach: First, the STAIR tool is used to rapidly review the literature and identify tasks that seem to have not received much attention, like accessing crime sites through a window, mapping/gathering evidence, and analyzing blood smears. Secondly, we present a prototype of a small drone that implements these three tasks with 75%, 85%, and 80% performance, to perform a minimal analysis of an indoor crime scene. Lessons learned are reported, toward guiding next work in the area.
@@ I INTRODUCTION
Automation of crime scene analysis (CSA) is an important problem: For example, violence alone was estimated to cost 14.76 USD trillion globally in 2017 [1]. When crimes occur behind closed doors, investigation can be required to reveal the truth of what transpired. Currently, however, investigators can suffer physical danger from criminals or traps at the scene, illnesses from contagious or toxic materials, and psychological harm [2]. Victims as well often do not receive justice, given numerous challenges such as human error, bias, contamination, understaffing and underfunding; degradation of evidence over time; and old-fashioned methods that can be inefficient and inaccurate such as measuring bloodstains by hand, paper and pencil drawings, carrying and changing light box filters, or working in the dark with flashlights and potentially missing evidence [2].111universalclass.com/articles/law/setting-crime-scene-perimeters.htm As in Fig. 1, we imagine that small drones like in Table I, that can navigate in tight indoor spaces, could help to reduce risks to investigators and facilitate justice for victims by accessing crime scenes quickly, reducing risks of contamination, and using sensory and visualizing modalities not available to humans to better capture and share information. Thereby, the idea is not that humans should be replaced, but that a drone can help humans by providing rapid initial processing.
Figure 1: Basic concept: a drone could help to quickly and safely analyze crime scenes, (a) accessing the scene to gain situation awareness, (b) gathering evidence, (c) and conducting initial analysis for an investigation.
TABLE I: Some nano drones that could be used for indoor crime scene analysis (CSA).
@@
AeroVironment Nano Hummingbird 1̃9 g
Blade Nano QX 22 g
Crazyflie 27 g
Zano 55 g
Teledyne FLIR Black Hornet 3/4 32 g/70 g
@@
DJI Ryze Tello 80 g
Here, the term "drone" is defined loosely to comprise unmanned aerial vehicles (UAV), unmanned aircraft systems (UAS), unmanned aircraft vehicle systems (UAVS), remotely piloted aerial vehicles (RPAV), and remotely piloted aircraft systems (RPAS), including quadcopters, quadrocopters, quadrotors, and micro aerial vehicles (MAVs). Furthermore, we follow the convention that "nano drones" are defined as less than 250 g.222euro-sd.com/2024/04/articles/37409/nano-uav-and-micro-uav-developments
A challenge is that it’s unclear which tasks nano drones could do, since CSA is highly complex. To gain insight, we follow a speculative prototyping approach [5], guided by the STAIR tool from forensic science, which states that an analyst must understand the Situation, carry out Tasks, Analyze evidence, and Investigate, to obtain Results [6].  [^3] https://online.campbellsville.edu/infographics/7-steps-of-a-crime-scene-investigation/ Thus, our contribution here is two-fold:
  • Theoretical. We speculate on how drones could be useful for crime scene analysis, including a list of some potential desired capabilities.
  • Practical. We prototype drones that can semi-autonomously carry out some actions within a mock-up crime scene.
The remainder of the paper is structured as follows: In Section II, we position the current paper in regard to previous work, identifying gaps. In Section III, we explore three gaps via prototyping, with evaluation results discussed in Section IV. Our aim is to provide insights that could inspire discussion and interest, toward bringing justice in a faster, more accurate way.
@@ II RELATED WORK
To survey the area, top results were analyzed from ACM, IEEE Xplore, and Google Scholar with the search phrase "(drone OR UAV OR robot) AND (crime)". Bibliographies of relevant papers were scoured, and some known papers also added. To try to avoid missing important information, we also followed up with some extra queries to Scite.AI and ChatGPT. Results outside of the scope of the paper were removed, regarding crimes by drones, security of drones, and forensic analysis of drones to catch human criminals. This led to 32 papers being reviewed (newest: 2024, oldest: 1991), which are summarized below according to the STAIR framework.
@@ II-A Situation Awareness
Previous work has begun to describe how robots can be piloted to go quickly to crime scenes (e.g. from rooftops in Sweden, like a mini-helicopter),444youtube.com/watch?v=wV8rNjvqM9A detect doors and windows,555github.com/sayedmohamedscu/YOLOv8-Door-detection-for-visually-impaired-people and enter buildings by opening doors,666youtube.com/watch?v=wV8rNjvqM9A, breaking walls or being thrown inside by a human.777wired.com/2016/07/11-police-robots-patrolling-around-world Drones can even acrobatically pass through narrow gaps [7]. Some drones capable of threat detection have also been designed to detect dangerous traps, pathogens, and criminals, such as Improvised Explosive Devices (IEDs) with radar, disease-carrying animal carcasses via thermal camera [8], and escaping or hiding people [9].888straitstimes.com/singapore/courts-crime/keeping-watch-from-the-skies-police-unveil-two-new-drones-for-crowd-management-search-and-rescue In regard to detecting offenses, cameras have also been used to recognize fighting [10]. More generally, robots have used observation to automatically infer rules underlying human behavior and interact, in games like chasing, follow-the-leader, and tag [11].
Thus, much remains to be done in this area. Our brainstorming suggested that gaps could include:
@@
  • Threat detection. Indoor hiding criminals and victims who are paused or moving to avoid detection (e.g. in a closet, under furniture, or behind a wall in a neighboring room) could be detected via radar or thermal cameras.
  • Offense detection. Foundation models and Large language models (LLMs) could be shown multimodal data (e.g., surveillance footage of crimes) and queried; i.e., to determine roles (even for groups coming and going), infer past/future actions, estimate levels of force, pinpoint focal points of a crime (location and cause, like where a person was stabbed or a fire or explosion started), detect anomalies (such as broken lamp), and generate crime scene sketches, etc.
@@ II-B Tasks
Robots have been designed to deal with threats like IEDs (improvised explosive devices) [12], disinfect complex surfaces with ultraviolet irradiation [13], and safely extract samples from corpses  [14]. To avoid psychological harm to humans or privacy infringements, video can be altered via blurring, masking, or mosaics [15]. Our previous work also lists various robots carrying weapons such as explosives, stun guns, pepper spray, water, firearms, or nets, some of which have been used to detect, negotiate with, or neutralize threats, including via computer vision [16]. A makeshift perimeter to protect evidence could also be created by dropping bricks at appropriate locations, a strategy which has been used previously by drones to construct a building [17].
As shown in Table II, previous exploratory work in securing evidence can be roughly split into work manually piloting a drone or using a lidar to take 3D scans of a mock-up crime scene to assess potential benefits or demerits of technology for CSA; variables of interest have mainly been bloodstains, guns, knives, and bodies. Furthermore, some datasets have been created that could be used to train a visual system to detect evidence for crime scene analysis, as exemplified in Table III. Various photogrammetry tools also exist that could be used to create 3D models, such as 3DF Zephyr, [^9] https://www.3dflow.net/3df-zephyr-photogrammetry-software/ Scaniverse, [^10] https://scaniverse.com/ Luma AI, [^11] https://lumalabs.ai/dream-machine OpenDroneMap, [^12] https://www.opendronemap.org/ and DroneDeploy. [^13] https://www.dronedeploy.com/
TABLE II: Crime Scene Evidence.
Urbanova et al. [18] drone/CV (3D scanning outdoors via a piloted drone and photogrammetry (Agisoft PhotoScan)) dummy, bones, and artificial blood, in grass
Georgiou et al. [19] drone/CV (piloting a DJI Spark drone outdoors for accurate, fast, reliable detection) colored square foam pads
Cooney et al. drone/CV [20] drone/CV (piloting a Ryze Tello drone indoors to detect via YOLO/heat traces) hidden cameras
Bucknell and Bassindale [21] drone (piloted a Parrot AR.Drone 2.0 at various heights to check the effects of downwash on sensitive evidence) textile fibres on various substrates
Rymansaib et al. [22] drone/sonar mannequin "body" (future target possibly also weapons, narcotics, or IEDs)
Araujo et al. [23] simulation of a drone detecting evidence based on AirSim/Unreal Engine/YOLO body, bloodstain, gun, and knife (augmented MS-COCO data set)
@@
Nandhini and Thinakaran [25] CV (classification 7-layer CNN) in infrared images firearms, knives, money, blood, animals, cars, and cellphones (some unclarity regarding datasets)
Buck et al. [26] 3D scanning bullet trajectories, and damage to bodies and objects
Esposito et al. [27] 3D scanning locations and distances between corpses, bloodstains, and weapons, bullet trajectories, paths where the victim might have moved before dying, balcony height
Galanakis et al. [28] 3D scanning "dead body" and various objects such as a phone, screws, tape, and a box (dataset released)
@@
TABLE III: Some datasets.
Weapons Weapon detection datasets141414dasci.es/transferencia/open-data/24705 9261, 10039, 23,672, 3000, 2078, 3255 images
Weapons Internet Movie Firearms Database151515imfdb.org 31,147 articles
Shoe prints Crime Scene Footwear Impressions Database [30] 170 samples
Shoe prints MUES-SR10KS2S dataset [31] 10,096 samples
Shoe prints FID-300 dataset [32] 1,475 samples (not public)
Various Roboflow Universe161616universe.roboflow.com 350 million+ images, 500,000+ datasets
Thus, much remains to be done in this area. Gaps include:
  • Handling threats. A drone could estimate backdrops and infer where and when force could be used; as well, drones could clear rooms and enter in a fast, effective, unpredictable, and safe way, like a SWAT team.
  • Geofencing. In relation to detecting where a crime took place, a drone could create a perimeter, establish a "path of contamination" or common approach path, and manage the scene–preventing unauthorized entry and improper actions.
  • Handling evidence. When mapping, useful information, like the distances between evidence (e.g., blood and body and weapon/objects) should also be calculated. Another important step after detecting evidence is gathering it (e.g., blood with a swab). For blood, presumptive testing could be conducted with luminol or tetramethylbenzidine (Hemastix) mounted to a drone. Other targets that could be important to detect include footprints/shoeprints, clothing (e.g., gloves, shoes), other weapons (bullets, bullet casings, clubs), restraints (e.g., cable ties, handcuffs), extra DNA evidence (skin, hairs, body fluids other than blood–e.g., like saliva on cigarettes, drinks, food, or cutlery), and objects which might have broken due to an altercation (e.g., glass/pottery fragments, paints, particulates, and fragment patterns (“fractography”)). Furthermore any object close to evidence such as a corpse or blood might also be interesting. Basic inference could be conducted on corpses (identity, pose, age, gender, and estimated time of death (e.g., based on body temperature detected with thermal camera)). As well, a thermal camera could be used to estimate when a person last touched evidence like weapons or drugs if the drone arrives fast enough.
@@
  • Cleaning. A robot such as Roomba could be used in simple cases.171717www.irobot.com/en_US/roomba.html More complex cases might require removing all carpet, scrubbing floors, and more intensive cleaning. Such a robot could be designed for easy washing and sterilization.
@@ II-C Analysis and Investigation
In the Analysis step, theories can be drawn from the raw data about motives, opportunities, and means. Theories from multiple sources (e.g., witness testimonies and physical evidence) are then compared in the Investigation step to narrow down what likely happened.
Various work is starting to be done with blood pattern analysis (BPA). For example, a review from Weber and Lednev describes how time since deposition (TSD) can be estimated using CV via magenta values, brightness, or blood pool/crack ratio/colorimetric analysis, despite challenges due to effects of substrate, environment (temperature, humidity), and contamination [33]. Bergman et al. also describe using CV (CNN-based classification) to discriminate between passive drip vs. active spatter bloodstains [34]. As well, various mathematical models exist; for example, in a crime scene with a broken teapot, a model could be used to infer how the teapot broke [35].
Regarding investigation, one robot was designed to publish a report summarizing anomalies it detected along with interviews from nearby people [36]. More generally, logic programming languages like Prolog, and inference approaches such as Maximum Likelihood Estimation (MLE) or its Bayesian counterpart Maximum a posteriori inference (MAP) can also be used to narrow down possibilities.
@@
  • Transfer stains. In some cases, blood patterns could carry information about how people moved and so-called "second locations". Moreover, composite scenes comprising various kinds of patterns could be analyzed.
@@
  • Motive Inference. For example, a missing or open safe could indicate a robbery; conversely, a murder victim with a wallet could indicate an alternative motive. Also, a distinctive modus operandi could suggest the work of a serial criminal.
Thus, while much inspiration can be taken from the literature, we did not encounter an overview of how drones could be used for crime scene analysis; however, it appeared that many opportunities exist for development, some of which we explore in the next section.

### v2

@@ Abstract
Technologies such as robotics, Artificial Intelligence (AI), and Computer Vision (CV) can be applied to crime scene analysis (CSA) to help protect lives, facilitate justice, and deter crime, but an overview of the tasks that can be automated has been lacking. Here we follow a speculative prototyping approach: First, the STAIR tool is used to rapidly review the literature and identify tasks that seem to have not received much attention, like accessing crime scenes through a window, mapping/gathering evidence, and analyzing blood smears. Secondly, we present a prototype of a small drone that implements these three tasks with 75%, 85%, and 80% performance, to perform a minimal analysis of an indoor crime scene. Lessons learned are reported, toward guiding next work.
@@ I INTRODUCTION
Automation of crime scene analysis (CSA) is an important problem: For example, violence alone was estimated to cost 14.76 USD trillion globally in 2017 [1]. When crimes occur behind closed doors, investigations can be required. Currently, however, investigations can be hindered by danger, human error, decomposition, and use of old-fashioned methods [2]: Investigators can suffer physical danger from criminals or traps at the scene, illnesses from contagious or toxic materials, and psychological harm such as vicarious trauma [3, 4]. Victims as well often do not receive justice [5], due to various causes:
  • Human Causes include bias [6], errors and contamination [7], as well as understaffing and underfunding [5].
  • Natural Causes include degradation of evidence over time [8].
  • Old-fashioned Methods, which could introduce inefficiencies or inaccuracies, include measuring bloodstains by hand using using rulers, strings, and protractors [9], sketching using paper, pencils, erasers, rulers, and clipboards [10], or carrying powerful light sources and filters (which can weigh ten kilograms) to work in dark rooms (e.g.,  [11, 12]).
As in Fig. 1, we imagine that small drones which can navigate in tight indoor spaces could help investigators and victims by accessing crime scenes quickly, reducing risks of contamination, and using sensory and visualizing modalities not available to humans to better capture and share information. Thereby, the idea is not that humans should be replaced, but that a drone can help humans by providing rapid initial processing.
Figure 1: Basic concept: a drone could help to quickly and safely analyze crime scenes, (a) accessing the scene to gain situation awareness, (b) gathering evidence, and (c) conducting initial analysis for an investigation.
Here, the term "drone" is defined loosely to comprise unmanned aerial vehicles (UAV), unmanned aircraft systems (UAS), unmanned aircraft vehicle systems (UAVS), remotely piloted aerial vehicles (RPAV), and remotely piloted aircraft systems (RPAS), including quadcopters, quadrocopters, quadrotors, (ultra-lightweight/tiny) micro aerial vehicles (MAVs), and "pocket drones". Furthermore, we follow the convention that "nano drones" are less than 250 g [13]. Various nano drones have been created for research, commercial, or military use, that could be adapted for CSA, as shown in Table I.
TABLE I: Some nano drones that could be used for indoor CSA.
@@
Fünfiiber [15] 18 g
AeroVironment Nano Hummingbird [16] 1̃9 g
DelFly Explorer [17] 20 g
Blade Nano QX [18] 22 g
Crazyflie [19] 27 g
(Custom) Blade mCX2 Helicopter [20, 21] 30 g
Teledyne FLIR Black Hornet 3/4 [22, 23] 32 g/70 g
(Edge-FS Pocket drone) [24] 40 g
(46g quadrotor) [25] 46 g
@@
DJI Ryze Tello [27] 80 g
A challenge is that it’s unclear which tasks nano drones could do, since CSA is highly complex. To gain insight, we follow a speculative prototyping approach [28], guided by the STAIR "tool" from forensic science [6, 29]. Speculative prototyping seeks to rapidly, openly, creatively and critically craft glimpses from possible future scenarios, starting with the question: "What if drones could help with CSA?" The STAIR tool models how investigators think and structure their investigations in a systematic, organized manner, which is important given the complexity and multiplicity of processes to be carried out. It embarks from the idea that an analyst must understand the Situation, carry out Tasks, Analyze evidence, and Investigate, to obtain Results [6, 29]. Within this context, our contribution here is two-fold:
  • Theoretical. We speculate on how drones could be useful for CSA, including a list of some potential desired capabilities.
  • Practical. We prototype some capabilities for drones within a mock-up crime scene.
Below, in Section II, we position the paper in regard to previous work, identifying gaps. Then, in Section III, we explore three gaps via prototyping, with evaluation results discussed in Section IV. Our aim is to provide insights that could inspire discussion and interest, toward bringing justice in a faster, more accurate way.
@@ II RELATED WORK
To survey the area, top results were analyzed from ACM, IEEE Xplore, and Google Scholar with the search phrase "(drone OR UAV OR robot) AND (crime)". Bibliographies of relevant papers were scoured, and some known papers also added. To try to avoid missing important information, we also followed up with some extra queries to Scite.AI and ChatGPT. Results outside of the scope of the paper were removed, regarding crimes by drones, security of drones, and forensic analysis of drones to catch human criminals. This led to 32 papers being reviewed (newest: 2024, oldest: 1991), which are summarized below following the categorization scheme of the STAIR tool.
@@ II-A Situation Awareness
Since the first "Drone as First Responder" (DFR) program started in 2018 in Chula Vista, California, drones like the Matrice 300 RTK have been remotely piloted thousands of times (e.g., from rooftops, like a mini-helicopter) to rapidly respond to 911 calls and provide enhanced situation awareness about outdoor incidents such as car accidents, vandalism, overdoses, and noise disturbances [30]. (Such programs have now also spread to other cities in the US such as New Orleans and New York, as well as to other countries such as Sweden, indicating growing interest [31].) Robots can furthermore detect doors and windows [32], and enter buildings by opening doors [33], breaking walls or being thrown inside by a human [34]. Drones can even acrobatically pass through narrow gaps [35]. Some drones capable of threat detection have also been designed to detect dangerous traps, pathogens, and criminals, such as Improvised Explosive Devices (IEDs) with radar, disease-carrying animal carcasses via thermal camera [36], and escaping or hiding people [37, 38]. Offenses such as fighting have also been visually recognized [39]. More generally, robots have used observation to automatically infer the underlying meaning of human behaviors, in games like chasing, follow-the-leader, and tag [40].
However, much remains to be done in this area. Our brainstorming suggested that gaps could include:
@@
  • Threat detection. Criminals and victims hiding indoors (who might be paused or moving to avoid detection, e.g. in a closet, under furniture, or behind a wall in a neighboring room) could be detected via radar or thermal cameras. Microphones could also capture sounds from voices, gunshots, or breaking glass, to detect shooter locations or fleeing suspects.
  • Offense detection. Foundation models and Large language models (LLMs) could be shown multimodal data (e.g., surveillance footage of crimes) and queried; i.e., to determine roles (even for groups coming and going), infer past/future actions, estimate levels of force, pinpoint focal points of a crime (location and cause, like where a person was stabbed or a fire or explosion started), detect anomalies (such as a broken lamp), and generate crime scene sketches, etc.
@@ II-B Tasks
Robots have been designed to deal with threats like IEDs [41], disinfect complex surfaces with ultraviolet irradiation [42], and safely extract samples from corpses  [43]. To avoid psychological harm to humans or privacy infringements, video can be altered via blurring, masking, or mosaics [44]. Our previous work also lists various robots carrying weapons such as explosives, stun guns, pepper spray, water, firearms, or nets, some of which have been used to detect, negotiate with, or neutralize threats, including via computer vision [45]. A makeshift perimeter to protect evidence could also be created by dropping bricks at appropriate locations, a strategy which has been used previously by drones to construct a building [46].
As shown in Table II, previous work in securing evidence can be roughly split into manual piloting of a drone or use of a lidar to take 3D scans of a mock-up crime scene, toward assessing potential benefits or demerits of technology for CSA; variables of interest have mainly been mock-up bloodstains, guns, knives, and bodies.
TABLE II: Prior Work Related to Detecting and Scanning Evidence.
Urbanova et al. [47] Drone/CV (3D scanning outdoors via a piloted DJI Phantom 2 drone with a GoPro HERO 4 and photogrammetry with Agisoft PhotoScan) Target: Dummy, bones, and artificial blood, in grass (detected by humans)
Georgiou et al. [48] Drone/CV (piloting a DJI Spark drone outdoors for accurate, fast, reliable detection, picking colors with Matlab) Colored square foam pads
Cooney et al. [49] Drone/CV (piloting a DJI Ryze Tello drone indoors to detect via YOLO/heat traces) Hidden cameras
Bucknell and Bassindale [50] Drone (piloting a Parrot AR.Drone 2.0 at various heights to check the effects of downwash on sensitive evidence) Textile fibres on various substrates
Rymansaib et al. [51] Drone/sonar A mannequin "body" (future target possibly also weapons, narcotics, or IEDs)
Araujo et al. [52] Simulation of a drone detecting evidence based on AirSim/Unreal Engine/YOLO A body, bloodstain, gun, and knife (augmented MS-COCO data set)
@@
Nandhini and Thinakaran [54] CV (classification via a 7-layer CNN) in infrared images Firearms, knives, money, blood, animals, cars, and cellphones (although with some unclarity regarding datasets)
Buck et al. [55] 3D scanning Bullet trajectories and damage to bodies and objects
Esposito et al. [56] 3D scanning Locations and distances between corpses, bloodstains, and weapons; bullet trajectories; paths where the victim might have moved before dying; and balcony height
Galanakis et al. [57] 3D scanning A "dead body" and various objects such as a phone, screws, tape, and a box (dataset released)
@@
Here, too, much remains to be done:
  • Handling threats. A drone could estimate backdrops and infer where and when force could be used; as well, drones could clear rooms and enter in a fast, effective, unpredictable, and safe way, like SWAT teams.
  • Geofencing. After detecting where a crime took place, a drone could create a perimeter, establish a "path of contamination" or "common approach path", and manage the scene–preventing unauthorized entry and improper actions.
  • Handling evidence. When mapping, useful information, like the distances between samples (e.g., blood, bodies, and weapons) should also be calculated. Another important step after detecting evidence is gathering (e.g., blood with a swab). For blood, presumptive testing could be conducted with luminol or tetramethylbenzidine (Hemastix) mounted to a drone. Other important detection targets could include footprints/shoeprints, clothing (e.g., gloves, shoes), other weapons (e.g., bullets, bullet casings, gunshot/gunpowder residue, clubs), restraints (e.g., cable ties, handcuffs), extra DNA evidence (e.g., skin, hairs, body fluids other than blood–like saliva on cigarettes, drinks, food, or cutlery), and broken objects (e.g., glass/pottery fragments, paints, particulates, bullet holes, and fragment patterns (“fractography”)). Furthermore, any object close to evidence such as a corpse or blood might also be interesting. Basic inference could be conducted on corpses (e.g., identity, pose, age, gender, and estimated time of death (based on body temperature)). In addition to lidars, low-light/infrared/ultraviolet cameras, and gas sensors to detect bodies (e.g., via putrescine, cadaverine, methane, hydrogen sulphide, or ammonia), a thermal camera could help estimate when a person last touched evidence like weapons or drugs if a drone arrives fast enough.
@@
  • Cleaning. A robot such as Roomba could be used in simple cases [59]. More complex cases might require removing all carpet, scrubbing floors, and more intensive cleaning. Such a robot could be designed to be easy to wash and sterilize.
@@ II-C Analysis and Investigation
In the Analysis step, theories can be drawn from raw data about motives, opportunities, and means. Theories from multiple sources (e.g., witness testimonies and physical evidence) are then compared in the Investigation step to narrow down what likely happened.
Various work explores blood pattern analysis (BPA): For example, a review from Weber and Lednev describes how time since deposition (TSD) can be estimated using CV via magenta values, brightness, or blood pool/crack ratio, despite challenges due to effects of substrate, contamination, and environment (temperature and humidity) [60]. Of the three main kinds of bloodstains, some focus seemed to have been placed on the passive drips and active spatter, which have been, e.g., discriminated by Bergman et al. using Convolutional Neural Networks (CNNs) [61]. We could only find a few examples of transfer stains (smears and swipes) in one dataset, which included slow-motion videos of a person swiping clean or bloodied glass or cloth with a finger, cloth, paper towel, or toothbrush from below or from the side [62]. (Although insightful, extra work would be required to adapt such data for a drone intended to fly over evidence after a crime has been committed.)
As well, various mathematical models exist that could be used to infer how an object (such as a teapot) broke [63]. Robots could also be designed to publish a report summarizing detected anomalies, along with interviews from nearby people [64]. Furthermore, logic programming languages like Prolog, and inference approaches such as Maximum Likelihood Estimation (MLE) or its Bayesian counterpart Maximum a posteriori inference (MAP) could also be used to narrow down possibilities.
@@
  • Transfer stains. In some cases, blood patterns could carry information about how victims and criminals moved and so-called "second locations". Moreover, composite scenes comprising various kinds of patterns could be analyzed.
@@
  • Motive Inference. A missing or open safe could indicate financial motives–the opposite, for a murder victim with wallet intact. Also, a distinctive modus operandi could suggest the work of a serial criminal.
Thus, our review did not reveal a pre-existing overview of how drones could be used for CSA; however, it seemed that much inspiration could be taken from the literature, and that many opportunities exist for development, some of which we explore in the next section.